Suppose the ratio of desktop to laptop computers in a school is 3 to 4. Give three scenarios that would result in the school having this ratio of computers. For each scenario, write the ratio of desktop to laptop computers and then simplify it.
1
Expert's answer
2019-09-04T11:26:07-0400

1) desktop = 30, laptop = 40

30/40=(3*10)/(4*10)=3/4


2) desktop = 12, laptop = 16

12/16=(3*4)/(4*4)=3/4


3) desktop = 21, laptop = 28

21/28=(3*7)/(4*7)=3/4